OpenStudy (anonymous):

Question 3 (Multiple Choice Worth 2 points) (05.01) How many solutions will this system of equations have? y = 3.5x - 3.5 y = -3.5x + 3.5 No solution Infinite solutions One solution Two solutions Question 4 (Multiple Choice Worth 1 points) (05.01) The graph of a system of equations will intersect at more than 1 point. Always Sometimes Never

OpenStudy (anonymous):

For question 1, the answer is B because those set of equations form parallel lines, which have infinite solutions. And the answer to Question 4, is D. because they are parallel.

OpenStudy (anonymous):

Question 3 is infinite

OpenStudy (anonymous):

I figured that @AnyTipzWouldHelp

OpenStudy (anonymous):

Well then you had no need to ask :)

OpenStudy (anonymous):

Dude really dont be rude @AnyTipzWouldHelp

OpenStudy (anonymous):

I was just stating... I had no intentions of being rude, just implying that he's smart enough to handle it...

OpenStudy (anonymous):

really everybody needs some confirmation can you help me with more

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

Okay then..sorry its just that I get really mad when people make rude comments over the internet ya know? Its so immature and its just wrong..and Yes @Swagged.Out.Jay

OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 1 points) (05.01) An inconsistent system of equations is a system with __________. the same line parallel lines intersecting lines and lines that have the same equation intersecting lines and lines that have the same slope Question 6 (Multiple Choice Worth 2 points) (05.01) Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5 Question 7 (Multiple Choice Worth 1 points) (05.01) The graph of a system of equations with different slopes will have no solutions. Always Sometimes Never

OpenStudy (anonymous):

Question 5: the answer is parallel lines because they are the only type of lines that are inconsistent. Question 7: D. because they do have solutions.

OpenStudy (anonymous):

Sorry, I'm not sure about question 6.

OpenStudy (anonymous):

I think 7 is never

OpenStudy (anonymous):

help @crabby_patty

OpenStudy (anonymous):

Question 2 (Multiple Choice Worth 1 points) (05.01) Choose the correct graph of the given system of equations. y + 2x = -1 3y - x = 4 graph of two lines that intersect at the point (negative 1, 1) graph of two lines that intersect at the point (1, 1) graph of two parallel lines with positive slopes None of the above

OpenStudy (anonymous):

I dont know how to do this one..sory..any others?

OpenStudy (anonymous):

no but thanks